an open macroeconomic model for a hypothetical economy is represented as follows

Y=Co+Io+Go+Xo-Mo, M=mo+m1yd, C=co+c1yd, T=tY and Yd=Y-T

a) Show that equal change in tax ans government expenditure are expansionary to the economy

b) Derive the equilibrium level of savings in the economy
1
Expert's answer
2016-01-19T08:44:11-0500
The equal change in tax and government expenditure are expansionary for the economy, because if the government expenditure increases (Go to G1), the GDP will increase too, as Y= C0 +Io+Go+X0-M. So, the equal change in tax and government expenditure will have expansionary effect.
In the equilibrium savings are equal to investment, so in our case the equilibrium level of savings is S = Io.
Investment multiplier is simply the multiplier effect of an injection of investment into an economy.
The investment multiplier in our case will be mi = 1/(1 - c) = 1/(1 - c1).